OpenStudy (anonymous):

need help with writing an equation for this problem. m = 5, passing through P(0, 7)

OpenStudy (anonymous):

y=5m+7

OpenStudy (hoblos):

apply the formula y-y1=m(x-x1) y-7=5(x-0) y=5x+7

OpenStudy (anonymous):

my book sya to use point-slope form to write the equation of the line with given properties.

OpenStudy (hoblos):

yes... that what we gave you!!

OpenStudy (rulnick):

point-slope form: y-7 = 5(x-0) done :)

OpenStudy (anonymous):

ok not to disagree with you but in the back of my book it says that it should be 5x-y= -7

OpenStudy (hoblos):

it is the same answer but in a different form we wrote it in slope-intercept form while in the back of your book it is written in standard form

OpenStudy (anonymous):

actually you wrote it in point slope form:)

OpenStudy (anonymous):

ok then I should have asked what the genarl form was.

OpenStudy (anonymous):

Yes and that is what I needed but in general form

OpenStudy (anonymous):

Sorry I am very confused about this math

OpenStudy (anonymous):

ok so can someone so me how you come up with this equation please?

OpenStudy (rulnick):

point-slope is y-y0 = m (x-x0) ... slope m through (x0,y0) standard is Ax + By = C slope-intercept is y = mx + b (m is slope, b is y-intercept) You said point slope, so that's the one I gave you above: y-7 = 5(x-0) ... slope 5 through (0,7). OK?